OpenStudy (anonymous):

How do you solve 1/4x=12

OpenStudy (amistre64):

personally, I wouldnt, it looks good just the way it is :)

OpenStudy (amistre64):

or if I wanted to; multiply both sides by 4

OpenStudy (anonymous):

well you couldent do that because i really meant 1 over 4 not 1 divided by 4

OpenStudy (amistre64):

ahhhh....... is it: 1/(4x) ??? = 12

OpenStudy (amistre64):

...and 1 over 4 IS 1 divided by 4......

OpenStudy (anonymous):

no like 1 over 4 as a fraction, then times x equals 12

OpenStudy (amistre64):

when we take a pie and divide up amongst the relatives at thanksgiving, we are handing our fractions of the pie to each person....fractions ARE divisions

OpenStudy (amistre64):

multiply both sides by 4

OpenStudy (anonymous):

Ok i get it now sorry :(

OpenStudy (amistre64):

if you get it, then no need to be sorry :)

OpenStudy (anonymous):

:) umm 1 more question since you are timesing x by 1 over 4 wouldent you divide 4 by both sides of the equal sign

OpenStudy (amistre64):

if we divide both sides by 4 that would be possible, but pointless in solveing for x: (1/4)x/4 = 12/4 (1/16)x = 3 doesnt help towards our cause....

OpenStudy (amistre64):

remeber, any number "over" itself is equal to "1" 4/4 = 1 (4/4)x = x

OpenStudy (amistre64):

(4)(1/4)x = 12(4) (4/4)x = 48 x = 48